On January 1, 2020, Shamrock Corporation purchased 321 of the $1,000 face value, 10%, 10-year bonds of Walters Inc. The bonds mature on January 1, 2030, and pay interest annually beginning January 1, 2021. Shamrock purchased the bonds to yield 11%. How much did Shamrock pay for the bonds? (Round factor values to 5 decimal places, e.g. 1.25124 and final answer to 0 decimal places, e.g. 458,581.) Shamrock must pay for the bonds $enter a dollar amount should be paid for the bonds
Answer #1
Amount PV factor 11% Present value
Annual interest 32100 5.88923 189044
Principal 321000 0.35218 113050
Total 302094
Shamrock must pay for the bonds 302094
Workings:
Annual interest 32100 =321000*10%
PV factor 11%:
Annual interest 5.88923 =(1-(1.11)^-10)/0.11
Principal 0.35218 =1/1.11^10